# Trajectory-to-Action Pipeline (TAP): Automated Scenario Description Extraction for Autonomous Vehicle Behavior Comparison

### v1

@@ 4.1. Threshold Calculation Problem
The objective of threshold calculation is to find separation thresholds $\Theta$ that ensure samples with the same label are more similar to each other than to those with different labels. Automatically deriving these thresholds from the data enables our method to generalize across datasets. Let $D$ denote the set $D=\{D_{\omega},D_{a},D_{v}\}$ of distributions of 1 second sampled averages of yaw rate $\omega$, acceleration $a$, and velocity $v$ across the entire dataset, spanning all scenarios. For example, vehicle $V^{i}_{j}$ contributes the 1 second sampled average of its accelerations $\{\bar{a}^{i}_{j,1},\bar{a}^{i}_{j,2},\dots\}$ to the distribution $D_{a}$. The sampling interval of 1 second allows noise to be smoothed without losing key behavioral differences. The velocity distribution $D_{v}$ and yaw rate distribution $D_{\omega}$ are defined similarly. The separating thresholds $\Theta=\{\Theta_{\omega},\Theta_{a},\Theta_{v}\}$ are defined over their corresponding distributions $D_{\omega}$, $D_{a}$, and $D_{v}$, where $\Theta_{\omega}=\{\theta_{\omega,\text{str}},\theta_{\omega,\text{grad}},\theta_{\omega,\text{med}}\}$, $\Theta_{a}=\{\theta_{a,\text{dec}},\theta_{a,\text{acc}}\}$, and $\Theta_{v}=\{\theta_{v,\text{stop}},\theta_{v,\text{slow}},\theta_{v,\text{med}}\}$ are thresholds that realize the rule-based function $f_{\Theta}$. For each distribution, the corresponding thresholds separate the distribution into distinct partitions that map vehicle states to labels. E.g. acceleration thresholds $\Theta_{a}$ separate the acceleration distribution $D_{a}$ into partitions: Decelerate, Maintain Speed, and Accelerate. The full list of partitions and the logical rules are included in Appendix B. Each sample that falls into the same partition is given the same SDL label. The distance between two samples within a partition $p1,p2\in D_{\text{part}}$ is the absolute value of their difference $|p1-p2|$ as a measure of their similarity.
(2) $\mu_{\text{part}}=\frac{\sum_{p1,p2\in D_{\text{part}}}|p1-p2|}{{|D_{\text{part}}|\choose 2}}.$
[…]
(3) $J(\Theta)=\sum_{part1,part2\in D}||\mu_{\text{part1}}-\mu_{\text{part2}}||^{2}$
(4) $\Theta^{*}=\operatorname*{arg\,min}_{\Theta}J(\Theta)$
4.2. SDL Similarity Search Problem
(5) $S^{L}_{\text{ref}}=\{L^{i}_{j}:g(L_{\text{ref}},L^{i}_{j})<d_{\text{sim}}\}$
5. Methodology
The goal of the Trajectory-to-Action Pipeline (TAP) is to automate the extraction of Scenario Description Language (SDL) labels from raw, unlabeled vehicle trajectory data. The absence of ground truth labels prevents the direct use of supervised learning approaches. TAP extracts SDL labels offline, for post hoc analysis of vehicle behavior. Figure 2 provides an overview of the entire Trajectory-to-Action Pipeline. The pipeline consists of multiple stages, each refining the classification of vehicle actions. Below, we detail the different stages involved in TAP.
5.1. Trajectory-to-Action Pipeline
Figure 2. Trajectory-to-Action Pipeline (TAP) starts with acceleration, velocity, and yaw rate data as input. At each stage, it uses automata and rules to produce SDL labels at the level of refinement corresponding to that stage. The final stage of TAP produces action-level SDL labels for every vehicle.
(6) $\begin{cases}\omega_{t}>\theta_{\omega,\text{str}}\,;&L^{\text{Trace}}_{\text{lat}}=\text{Left}\\ -\theta_{\omega,\text{str}}<\omega_{t}<\theta_{\omega,\text{str}}\,;&L^{\text{Trace}}_{\text{lat}}=\text{Straight}\\ \omega_{t}<-\theta_{\omega,\text{str}}\,;&L^{\text{Trace}}_{\text{lat}}=\text{Right}\end{cases}$
Consider the example shown in Figure 3. This figure shows the yaw rate $\omega$, acceleration $a$, and velocity $v$ at each time step for an example trajectory. At $t=0.0$ seconds, the yaw rate satisfies the criteria for the trace-level SDL label Straight. At $t=1.0$ seconds, the yaw rate falls below the threshold $-\theta_{\omega,\text{str}}$. Therefore, the extracted label is Right Turn as the automaton transitions to the Right node. At $t=4.6$ seconds, the yaw rate rises back above the threshold to again be assigned a Straight SDL label. Similar to the trace lateral automaton, the acceleration automaton produces trace-level longitudinal SDL labels $L^{\text{Trace}}_{\text{long}}$. Figure 3 also shows how the acceleration automaton produces longitudinal labels from the acceleration $a$ for our example. At the end of stage 1, TAP has automatically extracted trace-level SDL labels $L^{\text{Trace}}$ from trajectory yaw rate $\omega$ and acceleration $a$.
Figure 3. Example trajectory of a vehicle making a right turn, with yaw rate $\omega$, acceleration $a$, and velocity $v$ analyzed by TAP. Lateral SDL labels (trace-level) are Right Turn and Straight, while longitudinal labels are Maintaining Speed and Accelerating. At stage 4, TAP refines the velocity data into SDL actions: Maintain Slow Speed, Accelerate Slow Speed, and Accelerate Medium Speed.
[…]
In summary, for each of the stages in TAP, the output corresponds to the ordered sequence of labels $L^{\text{X}}_{\text{lat}}$ and $L^{\text{X}}_{\text{long}}$ at increasing levels of refinement. These labels are all the information necessary to construct the SDL label desribed in Section 3: $L^{i}_{j}=\langle\{L_{\text{lat},1},L_{\text{lat},2},\dots\}$,$\{L_{\text{long},1},L_{\text{long},2},\dots\}\rangle$ that capture the behavior of vehicle $V^{i}_{j}$ in scenario $S_{i}$.
5.2. Optimizing SDL Extraction Thresholds with Cross-Entropy
The accuracy of the SDL labels generated by TAP depends on the separation thresholds $\Theta$ which are used in the transition rules of the different TAP automata. The objective then is to automatically determine thresholds $\Theta$ from the distributions $D_{\omega}$, $D_{a}$, and $D_{v}$. This was described in Section 4.2 as the minimization optimization problem for $J(\Theta)$ (Equation 4). The optimization problem is a search over the manifold of $J(\Theta)$ to find thresholds $\Theta^{*}$ that minimize the objective function. For e.g. $J(\Theta_{\omega})$ is a 3-dimensional manifold over the yaw rate thresholds $\Theta_{\omega}=\{\theta_{\omega,\text{str}},\theta_{\omega,\text{grad}},\theta_{\omega,\text{med}}\}\in\mathbb{R}^{3}$. If $J(\Theta)$ were a smooth function, this problem is trivially solved by computing the gradient $\nabla J(\Theta)$. However, in our case, $J(\Theta)$ is defined by a set of disjointed rules (Table 2), which in-turn depend on the thresholds. As a result, due to $J(\Theta)$ non-smooth nature it is not differentiable, making $\nabla J(\Theta)$ unavailable in closed-form.
Consequentially, we need to approximate the gradient ^$\hat{\nabla}J(\Theta)$ using a successive approximation-based method. Such methods perform gradient descent on the parameters $\Theta$ in an iterative manner, beginning with an initial guess $\Theta_{0}$. At each iteration $k$, the parameters are updated using $\Theta_{k+1}=\Theta_{k}+\eta\hat{\nabla}J(\Theta_{k})$ where $\eta$ is the learning rate, and ^$\hat{\nabla}J(\Theta_{k})$ is an approximation of the gradient of $J(\Theta_{k})$. Iterating using this successive approximation yields the optimal values of the thresholds, $\Theta^{*}$. TAP uses the cross-entropy method (CEMethod, ) for approximating the gradient ^$\hat{\nabla}J(\Theta_{k})$. The cross-entropy method approximates the gradient ^$\hat{\nabla}J(\Theta_{k})$ using finite differences:
(7) ^$\hat{\nabla}J(\Theta_{k})=\frac{J(\Theta_{k}+\epsilon)-J(\Theta_{k}-\epsilon)}{2\epsilon}$
6. Experiments and Results
6.1. Dataset and Preprocessing
We use the Waymo Open Motion Dataset (WOMD) (WOMD_2021, ) as our source of scenarios $S$. TAP is evaluated on a representative subset of 492 scenarios from WOMD, containing $25,889$ vehicle trajectories, each recorded at 10 Hz over a duration of $T=9.0$ seconds. Each scenario $S_{i}$ includes data from several vehicles $V^{i}_{j}$ including their states $\{x,y,v,\phi\}$, where $x,y$ denote the vehicle’s position in map coordinates, $v$ represents its velocity in m/s, and $\phi$ is the heading in radians. To obtain the complete state vector $s=\{x,y,v,a,\phi,\omega\}$, we derive acceleration $a$ and yaw rate $\omega$ from the recorded velocity and heading. Each vehicle’s trajectory is normalized to a common initial condition by translating positions $x,y$ to the origin and adjusting the heading to $\phi=0$ radians, enabling fair trajectory comparisons. Using the full state vector $s=\{x,y,v,a,\phi,\omega\}$, we construct 1-second average distributions $D_{a}$, $D_{v}$, and $D_{\omega}$ (Section 4.1), which serve as the foundation for TAP’s automatic threshold computation. WOMD trajectories do not contain ground truth action labels.
6.2. Result 1 - Automated Threshold Optimization
Figure 4. (a) Data distributions with initial seeds shown as dotted lines. (b) Initial subcategory averages for each seed. (c) Objective function $J(\Theta)$ convergence per epoch. (d) Final subcategory averages after optimization; some skew remains due to inherent data imbalances, but all seeds converge to the same thresholds
Hyperparameters and Ablation: The cross-entropy method requires initial threshold seeds $\Theta^{1}_{0},\Theta^{2}_{0},\Theta^{3}_{0}$ to ensure effective convergence. We selected these seeds by analyzing the data distribution as well as using domain knowledge to achieve a starting point guess at balanced dataset partitions for cross-entropy. To enhance robustness and avoid local minima, three distinct seeds were used for each threshold. This approach improves data separability and ensures more reliable SDL extraction. The hyperparameters are listed in Table 1. Figure 4(a) also shows these initial threshold guesses plotted for their respective distributions. These thresholds define initial intra-partition separation values (Figure 4(b)), $\mu_{part}$, which measure the distance between samples within each partition with higher values indicating greater separation.
Evaluation: The cross-entropy method aims to optimize these thresholds for balanced separation across partitions. Figure 4(c) illustrates the convergence of the objective function $J(\Theta)$ over multiple epochs, with similar final values for each seed indicating optimal thresholds. As $J(\Theta)$ decreases, it demonstrates effective learning by the method. Figure 4(d) shows the final intra-partition separations, where $\mu_{part}$ is more balanced across partitions compared to the initial values (Figure 4(b)). The lateral thresholds converged to values of $\Theta_{\omega}=\{0.0283,0.0754,0.1541\}$ rad/s. The longitudinal thresholds converged to $\Theta_{a}=\{-1.3715,1.5557\}$ m/s2 and $\Theta_{v}=\{0.1,10.2140,24.4046\}$ m/s. Consistent convergence across all seeds confirms the cross-entropy method’s reliability in optimizing partition balance. The complete table of initial seeds and converged values is included in Appendix C.
$\epsilon$ 0.05 0.05 0.005
$\eta$ 0.05 0.2 0.01
Table 1. The hyper-parameters $\epsilon$ and learning rate used for each category of the cross-entropy method.
6.3. Result 2 - SDL Similarity Search
We use TAP to extract SDL labels for each trajectory, as outlined in Section 5.1. The objective of SDL similarity search is to find trajectories with similar behaviors by measuring the distance between their respective SDL labels. We use the distance function $g_{\text{SDL}}(L_{\text{ref}},L^{i}_{j})=0$ if $L_{\text{ref}}=L^{i}_{j}$, and $g_{\text{SDL}}(L_{\text{ref}},L^{i}_{j})=1$ otherwise, which considers two trajectories similar if they share the same SDL label. Other distance metrics, such as Levenshtein distance, could also be used for comparing SDL labels. WOMD trajectories do not contain ground truth labels for similarity, therefore, we compare TAP’s SDL-labels against two commonly used trajectory similarity metrics, Average Displacement Error (ADE) (ADE_similarity1, ; ADE_similarity2, ) and Dynamic Time Warping (DTW) (DTW_similarity1, ; DTW_similarity2, ). These metrics are used to compare distance between trajectories in autonomous driving prediction, measuring similarity through the Euclidean distances between corresponding points along trajectories.
Figure 5. Trajectories with similar SDL labels to the reference (magenta). The trajectory with the lowest ADE is shown in dark blue, and lowest DTW in light blue. Red labels indicate extra actions, while underlined labels indicate missing actions. Vehicle velocity plots are shown on the right. Additional similarity examples are included in Appendix D.
Figure 5 demonstrates TAP’s effectiveness in identifying behaviorally similar trajectories compared to baseline metrics. Given a reference trajectory $S^{L}_{\text{ref}}$ (magenta), TAP retrieves all scenarios with matching SDL labels (gray). The most similar trajectories identified by ADE (dark blue) and DTW (light blue) are also shown, along with TAP’s SDL labels for those trajectories. In Figure 5 example 1, the lateral motion of the reference trajectory includes a gradual left turn (triangular marker) followed by straight (circular marker) and the longitudinal behavior is at medium speed. This is an example where trajectories returned by both ADE and DTW also match this behavior, yielding trajectories with the same SDL labels as TAP. However, examples 2 and 3 reveal cases where ADE and DTW fail to capture key behaviors. In 2, both baseline methods miss the turning behavior in the reference trajectory, while TAP identifies trajectories with similar behavior despite slight spatial differences. In 3, ADE and DTW match the lateral SDL label but fail to capture the longitudinal behavior in acceleration, highlighting TAP’s superior accuracy in comparing similar trajectories. For empirical comparison, we conducted similarity searches on 25,889 trajectories. ADE failed to retrieve behaviorally similar trajectories for 7,924 samples $(30.56\%)$, often selecting spatially close trajectories with differing behaviors. DTW exhibited similar issues, with 6,246 cases $(24.13\%)$ failing to match the reference’s behavior. These examples demonstrate the limitations of using traditional metrics for behavior-based similarity searches. ADE and DTW focus only on spatial proximity, which can miss critical driving behaviors. In contrast, TAP’s SDL-based approach excels in identifying behaviorally similar trajectories, enabling more reliable comparisons for robust AV testing and real-world scenario evaluation. While this result focuses on comparing one trajectory at a time, our ongoing work extends this to multiple trajectory comparisons, enabling scenario-level analysis for more comprehensive behavior evaluations.
6.4. Result 3 - Unique Behaviors
Figure 6. Two trajectories whose labels were completely unique within the dataset. Both have many lateral and longitudinal actions, which is uncommon among the dataset, resulting in unique samples.
7. Discussion and Conclusion
This paper introduced the Trajectory-to-Action Pipeline (TAP), a method for automatically extracting interpretable Scenario Description Language (SDL) labels from trajectory data. TAP employs rule-based automata with data-derived thresholds, optimized using cross-entropy to generate structured, hierarchical labels. Using the Waymo Open Motion Dataset for validation, TAP demonstrated a 30% improvement in precision over Average Displacement Error (ADE) and a 24% improvement over Dynamic Time Warping (DTW) in identifying behaviorally similar trajectories. TAP also enables the automatic detection of unique scenarios within large datasets. Additionally, TAP enables the automatic detection of unique scenarios within a large dataset. Despite these contributions, there are key areas for future improvements. Currently, TAP focuses solely on vehicle trajectories, without integrating static environmental elements (e.g., road geometry) or accounting for interactions between multiple agents. Future work would enhance TAP by incorporating these to support more complex scenario comparisons. Additionally, while TAP introduces its own SDL schema, adapting the method to established frameworks like Scenic or OpenScenario could improve interoperability with existing testing and simulation tools. Lastly, TAP could be augmented using recent transformer-based architectures for encoding trajectories. Ultimately, TAP lays a strong foundation for automating behavior-based analysis in AV testing and safety evaluation, with potential for significant extensions to better reflect real-world scenario complexity.
Appendix
Appendix A Code Release and Reproducibility
We prioritize transparency and reproducibility by providing access to the methods and data used in this work. The Waymo Open Motion Dataset employed in our experiments is publicly available at https://waymo.com/open/. The implementation of the Trajectory-to-Action Pipeline (TAP), which automates the extraction of SDL embeddings, will be released on GitHub upon the paper’s publication. This release will include all necessary scripts and documentation to facilitate replication and further exploration of our approach.
Appendix B Partition Rules
In order to automatically extract SDL labels from trajectories, the vehicle state data distributions $D_{\omega}$, $D_{a}$, and $D_{v}$ are separated by on a set of thresholds $\Theta=\{\Theta_{\omega},\Theta_{a},\Theta_{v}\}$. These thresholds separate the yaw rate, acceleration, and velocity of the vehicle into distinct partitions corresponding to the SDL Labels. Table 2 shows the partitions and the rules-based criteria for each data distribution. For example, the yaw rate distribution is partitioned by the thresholds $\Theta_{\omega}=\{\theta_{\omega,\text{str}},\theta_{\omega,\text{grad}},\theta_{\omega,\text{med}}\}$ into labels of Straight, Gradual Turn, Medium Turn, and Aggressive Turn.
@@ Appendix B Partition Rules
Table 2. The rule-based criteria for data partitions. For instance, vehicle acceleration is separated by $\theta_{a,\text{acc}}$ and $\theta_{a,\text{dec}}$ into partitions of Decelerate, Maintain Speed, and Accelerate.
Appendix C Cross Entropy Seeds
Cross-entropy optimization can risk converging to local minima. To mitigate this, we initialized the optimization process with three different threshold guesses for each data distribution. Each guess was run for 100 epochs, using the hyperparameters listed in Table 1. Despite the variations in initial thresholds, the method consistently converged to similar final values across all runs.
Tables 3 to 5 present the initial threshold guesses and the corresponding final values for acceleration, velocity, and yaw rate. The consistency of convergence suggests that the final thresholds provide well-balanced partitions and that the optimization process successfully avoids local minima. TAP adopts the thresholds that yield the lowest objective function value as its final parameters.
Threshold ($\Theta_{a}$) Initial Value ($m/s^{2})$ Final Value ($m/s^{2}$)
Seed 1
$\theta_{a,\text{dec}}$0.5 -1.2350
$\theta_{a,\text{acc}}$ 0.5 1.4870
Seed 2
$\theta_{a,\text{dec}}$1.0 -1.3196
$\theta_{a,\text{acc}}$ 1.0 1.5379
Seed 3
$\theta_{a,\text{dec}}$2.0 -1.3715
$\theta_{a,\text{acc}}$ 2.0 1.5557
Table 3. Initial seeds and final optimized acceleration thresholds. The final thresholds with the lowest objective function (shown in bold) were used for TAP.
Threshold ($\Theta_{v}$) Initial Value ($m/s$) Final Value ($m/s$)
Seed 1
$\theta_{v,\text{slow}}$ 7.0 10.2140
$\theta_{v,\text{med}}$ 22.0 24.4046
Seed 2
$\theta_{v,\text{slow}}$ 10.0 10.2098
$\theta_{v,\text{med}}$ 24.0 24.3755
Seed 3
$\theta_{v,\text{slow}}$ 12.0 10.2099
$\theta_{v,\text{med}}$ 26.0 24.3568
Table 4. Initial seeds and final optimized velocity thresholds. The final thresholds with the lowest objective function (shown in bold) were used for TAP.
Threshold ($\Theta_{\omega}$) Initial Value ($rad/s$) Final Value ($rad/s$)
Seed 1
$\theta_{\omega,\text{str}}$ 0.01 0.0273
$\theta_{\omega,\text{grad}}$ 0.05 0.0728
$\theta_{\omega,\text{med}}$ 0.1 0.1517
Seed 2
$\theta_{\omega,\text{str}}$ 0.02 0.0283
$\theta_{\omega,\text{grad}}$ 0.07 0.0754
$\theta_{\omega,\text{med}}$ 0.15 0.1541
Seed 3
$\theta_{\omega,\text{str}}$ 0.05 0.0320
$\theta_{\omega,\text{grad}}$ 0.1 0.0898
$\theta_{\omega,\text{med}}$ 0.2 0.1585
Table 5. Initial seeds and final optimized yaw rate thresholds. The final thresholds with the lowest objective function (shown in bold) were used for TAP.
Appendix D Additional Similarity Results
In this section, we provide further examples demonstrating TAP’s ability to identify behaviorally similar trajectories more effectively than the baseline metrics, Average Displacement Error (ADE) and Dynamic Time Warping (DTW). Figure 7 presents a reference trajectory (magenta), along with trajectories that match the extracted SDL labels (gray) and those identified as most similar by ADE (dark blue) and DTW (light blue). Each trajectory’s SDL label, as extracted by TAP, is also shown.
Example 1
In example 1 of Figure 7, the reference trajectory travels straight, makes an aggressive right turn, and then straightens again. The nearest ADE trajectory starts straight and makes a similar aggressive right turn but fails to straighten afterward. In contrast, the nearest DTW trajectory begins its turn prematurely, missing the initial straight phase entirely. Additionally, both baseline trajectories fail to replicate the reference’s longitudinal behavior: the reference begins at a slow speed, accelerates to a medium speed, and then maintains that speed. Both ADE and DTW trajectories continue accelerating beyond the medium speed.
Example 2
In example 2, the reference trajectory travels straight before executing an aggressive left turn and accelerating slightly before stabilizing at a slow speed. The nearest ADE and DTW trajectories replicate the lateral behavior but deviate significantly in longitudinal dynamics. The ADE trajectory begins at a medium speed and decelerates to a slow speed, while the DTW trajectory fails to accelerate enough to be classified as accelerating.
Example 3
In example 3, the reference trajectory makes a gradual left turn between two straight segments, all at a medium speed. The nearest ADE trajectory matches the lateral pattern but travels at a much slower speed. Conversely, the nearest DTW trajectory mirrors the longitudinal behavior but transforms the entire sequence into a prolonged turn without straight-line segments.
These examples highlight how TAP more accurately captures both lateral and longitudinal dynamics, leading to better identification of behaviorally similar trajectories compared to baseline methods.

### v2

@@ IV-A Threshold Calculation Problem
The objective of threshold calculation is to find separation thresholds $\Theta$ that ensure samples with the same label are more similar to each other than to those with different labels. Automatically deriving these thresholds from the data enables our method to generalize across datasets. Let $D$ denote the set $D=\{D_{\omega},D_{a},D_{v}\}$ of distributions of 1 second sampled averages of yaw rate $\omega$, acceleration $a$, and velocity $v$ across the entire dataset, spanning all scenarios. For example, vehicle $V^{i}_{j}$ contributes the 1 second sampled average of its accelerations $\{\bar{a}^{i}_{j,1},\bar{a}^{i}_{j,2},\dots\}$ to the distribution $D_{a}$. The sampling interval of 1 second allows noise to be smoothed without losing key behavioral differences. The velocity distribution $D_{v}$ and yaw rate distribution $D_{\omega}$ are defined similarly. The separating thresholds $\Theta=\{\Theta_{\omega},\Theta_{a},\Theta_{v}\}$ are defined over their corresponding distributions $D_{\omega}$, $D_{a}$, and $D_{v}$, where $\Theta_{\omega}=\{\theta_{\omega,\text{str}},\theta_{\omega,\text{grad}},\theta_{\omega,\text{med}}\}$, $\Theta_{a}=\{\theta_{a,\text{dec}},\theta_{a,\text{acc}}\}$, and $\Theta_{v}=\{\theta_{v,\text{stop}},\theta_{v,\text{slow}},\theta_{v,\text{med}}\}$ are thresholds that realize the rule-based function $f_{\Theta}$. For each distribution, the corresponding thresholds separate the distribution into distinct partitions that map vehicle states to labels. E.g. acceleration thresholds $\Theta_{a}$ separate the acceleration distribution $D_{a}$ into partitions: Decelerate, Maintain Speed, and Accelerate. The full list of partitions and the logical rules are included in Table I. Each sample that falls into the same partition is given the same SDL label. The distance between two samples within a partition $p1,p2\in D_{\text{part}}$ is the absolute value of their difference |p1p2||p1-p2| as a measure of their similarity.
@@
TABLE I: The rule-based criteria for data partitions. For instance, vehicle acceleration is separated by $\theta_{a,\text{acc}}$ and $\theta_{a,\text{dec}}$ into partitions of Decelerate, Maintain Speed, and Accelerate.
$\mu_{\text{part}}=\frac{\sum_{p1,p2\in D_{\text{part}}}|p1-p2|}{{|D_{\text{part}}|\choose 2}}$. (2)
This value is calculated as the sum of all pairwise distances between samples within the partition, divided by the total number of sample pairs. A lower value for $\mu_{\text{part}}$ indicates higher similarity among the samples in that partition. Given two partitions $\text{part1},\text{part2}\in D$ we can now define the inter-partition similarity as $\mu_{\text{part1}}-\mu_{\text{part2}}$. Therefore, objective function
[…]
$\Theta^{*}=\operatorname*{arg\,min}_{\Theta}J(\Theta)$ (4)
IV-B SDL Similarity Search Problem
$S^{L}_{\text{ref}}=\{L^{i}_{j}:g(L_{\text{ref}},L^{i}_{j})<d_{\text{sim}}\}$ (5)
V Methodology
Figure 2: Trajectory-to-Action Pipeline (TAP) starts with acceleration, velocity, and yaw rate data as input. At each stage, it uses automata and rules to produce SDL labels at the level of refinement corresponding to that stage. The final stage of TAP produces action-level SDL labels for every vehicle.
The goal of the Trajectory-to-Action Pipeline (TAP) is to automate the extraction of Scenario Description Language (SDL) labels from raw, unlabeled vehicle trajectory data. The absence of ground truth labels prevents the direct use of supervised learning approaches. TAP extracts SDL labels offline, for post hoc analysis of vehicle behavior. Figure 2 provides an overview of the entire Trajectory-to-Action Pipeline. The pipeline consists of multiple stages, each refining the classification of vehicle actions. The implementation of TAP will be released on GitHub upon the paper’s publication. Below, we detail the different stages involved in TAP.
V-A Trajectory-to-Action Pipeline
$\begin{cases}\omega_{t}>\theta_{\omega,\text{str}}\,;&L^{\text{Trace}}_{\text{lat}}=\text{Left}\\ -\theta_{\omega,\text{str}}<\omega_{t}<\theta_{\omega,\text{str}}\,;&L^{\text{Trace}}_{\text{lat}}=\text{Straight}\\ \omega_{t}<-\theta_{\omega,\text{str}}\,;&L^{\text{Trace}}_{\text{lat}}=\text{Right}\end{cases}$ (6)
Consider the example shown in Figure 3. This figure shows the yaw rate $\omega$, acceleration $a$, and velocity $v$ at each time step for an example trajectory. At $t=0.0$ seconds, the yaw rate satisfies the criteria for the trace-level SDL label Straight. At $t=1.0$ seconds, the yaw rate falls below the threshold $\theta_{\omega,\text{str}}$. Therefore, the extracted label is Right Turn as the automaton transitions to the Right node. At $t=4.6$ seconds, the yaw rate rises back above the threshold to again be assigned a Straight SDL label. Similar to the trace lateral automaton, the acceleration automaton produces trace-level longitudinal SDL labels $L^{\text{Trace}}_{\text{long}}$. Figure 3 also shows how the acceleration automaton produces longitudinal labels from the acceleration $a$ for our example. At the end of stage 1, TAP has automatically extracted trace-level SDL labels $L^{\text{Trace}}$ from trajectory yaw rate $\omega$ and acceleration $a$.
Figure 3: Example trajectory of a vehicle making a right turn, with yaw rate $\omega$, acceleration $a$, and velocity $v$ analyzed by TAP. Lateral SDL labels (trace-level) are Right Turn and Straight, while longitudinal labels are Maintaining Speed and Accelerating. At stage 4, TAP refines the velocity data into SDL actions: Maintain Slow Speed, Accelerate Slow Speed, and Accelerate Medium Speed.
Stage 4 - Action At the beginning of stage 4, every trajectory $\tau^{j}$ has maneuver-level SDL labels $L^{\text{Maneuver}}_{\text{lat}}=\{\text{Left Turn},\allowbreak\text{Left Merge},\allowbreak\text{Straight},\allowbreak\text{Right Turn},\allowbreak\text{Right Merge}\}$ and $L^{\text{Maneuver}}_{\text{long}}=\{\text{Accelerate},\allowbreak\text{Stopped},\allowbreak\text{Maintain Speed},\allowbreak\text{Decelerate}\}$. In the Action stage, the goal is to distill the maneuver-level SDL labels into more refined action-level SDL labels $L^{\text{Action}}$ that can capture even more nuanced lateral and longitudinal behaviors. Not all turns are equal—some are gradual, while others are sharp. To capture this, action-level SDL labels classify turns as Gradual Turn, Medium Turn, and Aggressive Turn. Similarly, longitudinal actions like Accelerate, Maintain Speed, and Decelerate are further divided into speed profiles: Slow, Medium, and Fast. To refine lateral maneuvers, TAP uses the action lateral automaton with thresholds $\theta_{\omega,\text{str}}$, $\theta_{\omega,\text{grad}}$, and $\theta_{\omega,\text{med}}$ as defined in Table I. This automaton refines the maneuver-level labels into action-level SDL labels such as Gradual, Medium, and Aggressive Turns. Given a turn maneuver $L^{\text{Maneuver}}_{\text{lat},t}$ and yaw rate $\omega_{t}$, the output is the action-level label $L^{\text{Action}}_{\text{lat},t}$. During this stage, TAP also uses the velocity profile automaton which has nodes of Slow, Medium, and Fast. The input to the velocity profile automaton is a longitudinal maneuver $L^{\text{Maneuver}}_{\text{long},t}$ and velocity $v_{t}$, and it assigns one of three different velocity profile labels to produce the action-level longitudinal SDL label $L^{\text{Action}}_{\text{long},t}$. Referring back to Figure 3, the bottom plot shows the vehicle’s velocity. It starts at a Slow speed (14 mph) at $t=0.0$ seconds and reaches Medium speed (22 mph) by $t=5.6$ seconds. In the previous TAP stage, two longitudinal maneuvers - Maintain Speed and Accelerate - were identified. The velocity profile automaton further refines these into Maintain Slow Speed, Accelerate Slow Speed, and Accelerate Medium Speed. The original maneuvers can sometimes be split into action-level labels shorter than 1 second. This poses a risk of data loss, as our SDL requires actions to be at least 1 second long. To address this, TAP assigns a single action based on the average value of the maneuver when the resulting actions are too short. After the action stage, every trajectory $\tau^{j}$ has been labeled with the lateral actions $L^{\text{Action}}_{\text{lat}}$ and the longitudinal actions $L^{\text{Action}}_{\text{long}}$ which are the most refined SDL labels.
In summary, for each of the stages in TAP, the output corresponds to the ordered sequence of labels $L^{\text{X}}_{\text{lat}}$ and $L^{\text{X}}_{\text{long}}$ at increasing levels of refinement. These labels are all the information necessary to construct the SDL label desribed in Section III: $L^{i}_{j}=\langle\{L_{\text{lat},1},L_{\text{lat},2},\dots\}$$\{L_{\text{long},1},L_{\text{long},2},\dots\}\rangle$ that capture the behavior of vehicle $V^{i}_{j}$ in scenario $S_{i}$.
V-B Optimizing SDL Extraction Thresholds with Cross-Entropy
The accuracy of the SDL labels generated by TAP depends on the separation thresholds $\Theta$ which are used in the transition rules of the different TAP automata. The objective then is to automatically determine thresholds $\Theta$ from the distributions $D_{\omega}$, $D_{a}$, and $D_{v}$. This was described in Section IV-B as the minimization optimization problem for $J(\Theta)$ (Equation 4). The optimization problem is a search over the manifold of $J(\Theta)$ to find thresholds $\Theta^{*}$ that minimize the objective function. For e.g. $J(\Theta_{\omega})$ is a 3-dimensional manifold over the yaw rate thresholds $\Theta_{\omega}=\{\theta_{\omega,\text{str}},\theta_{\omega,\text{grad}},\theta_{\omega,\text{med}}\}\in\mathbb{R}^{3}$. If $J(\Theta)$ were a smooth function, this problem is trivially solved by computing the gradient $\nabla J(\Theta)$. However, in our case, $J(\Theta)$ is defined by a set of disjointed rules (Table I), which in-turn depend on the thresholds. As a result, due to $J(\Theta)$ non-smooth nature it is not differentiable, making $\nabla J(\Theta)$ unavailable in closed-form.
Figure 4: (a) Data distributions with initial seeds shown as dotted lines. (b) Initial subcategory averages for each seed. (c) Objective function $J(\Theta)$ convergence per epoch. (d) Final subcategory averages after optimization; some skew remains due to inherent data imbalances, but all seeds converge to the same thresholds
Consequentially, we need to approximate the gradient $\hat{\nabla}J(\Theta)$ using a successive approximation-based method. Such methods perform gradient descent on the parameters $\Theta$ in an iterative manner, beginning with an initial guess $\Theta_{0}$. At each iteration $k$, the parameters are updated using $\Theta_{k+1}=\Theta_{k}+\eta\hat{\nabla}J(\Theta_{k})$ where $\eta$ is the learning rate, and $\hat{\nabla}J(\Theta_{k})$ is an approximation of the gradient of $J(\Theta_{k})$. Iterating using this successive approximation yields the optimal values of the thresholds, $\Theta^{*}$. TAP uses the cross-entropy method [31] for approximating the gradient $\hat{\nabla}J(\Theta_{k})$. The cross-entropy method approximates the gradient $\hat{\nabla}J(\Theta_{k})$ using finite differences:
$\hat{\nabla}J(\Theta_{k})=\frac{J(\Theta_{k}+\epsilon)-J(\Theta_{k}-\epsilon)}{2\epsilon}$ (7)
VI Experiments and Results
VI-A Dataset and Preprocessing
We use the publicly available Waymo Open Motion Dataset (WOMD) [32] as our source of scenarios $S$. TAP is evaluated on a representative subset of 492 scenarios from WOMD, containing $25,889$ vehicle trajectories, each recorded at 10 Hz over a duration of $T=9.0$ seconds. Each scenario $S_{i}$ includes data from several vehicles $V^{i}_{j}$ including their states $\{x,y,v,\phi\}$, where $x,y$ denote the vehicle’s position in map coordinates, $v$ represents its velocity in m/s, and $\phi$ is the heading in radians. To obtain the complete state vector $s=\{x,y,v,a,\phi,\omega\}$, we derive acceleration $a$ and yaw rate $\omega$ from the recorded velocity and heading. Each vehicle’s trajectory is normalized to a common initial condition by translating positions $x,y$ to the origin and adjusting the heading to $\phi=0$ radians, enabling fair trajectory comparisons. Using the full state vector $s=\{x,y,v,a,\phi,\omega\}$, we construct 1-second average distributions $D_{a}$, $D_{v}$, and $D_{\omega}$ (Section IV-A), which serve as the foundation for TAP’s automatic threshold computation. WOMD trajectories do not contain ground truth action labels.
VI-B Result 1 - Automated Threshold Optimization